40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 19 23:59:22 2020, anchal, Summary, ALS, ALS noise budget update ALS_nb_ExtraPoles.pdfALS_controls.yamlALS_nb_Kiwamus_Values.pdfALS_controls_Kiwamus_Values.yamlALS_simulink_model.svg
    Reply  Sun Sep 20 23:12:13 2020, rana, Summary, ALS, ALS noise budget update 
       Reply  Tue Sep 22 00:14:43 2020, anchal, Summary, ALS, ALS noise budget update ALS_NoiseBudgetUpdate.pdfALS_controls.yaml
          Reply  Tue Sep 22 12:14:42 2020, rana, Summary, ALS, ALS noise budget update 
             Reply  Wed Sep 23 11:13:49 2020, anchal, Summary, ALS, ALS noise budget update ALS_NoiseBudgetUpdate.pdfALS_controls.yaml
                Reply  Wed Oct 7 16:56:23 2020, anchal, Summary, ALS, ALS noise budget update - Updated AUX PDH Loop values Extract_X_AUX_PDH_Model.zipALS_NoiseBudgetUpdate.pdf
                   Reply  Thu Oct 8 11:59:52 2020, rana, Summary, ALS, ALS noise budget update - Updated AUX PDH Loop values 
                      Reply  Fri Oct 9 18:32:14 2020, anchal, Summary, ALS, ALS noise budget update - Updated AUX PDH Loop values 
                         Reply  Wed Oct 14 17:03:55 2020, anchal, Summary, ALS, ALS noise budget update - Added whitening filter for ADC ALS_NoiseBudgetUpdate.pdf
Message ID: 15622     Entry time: Fri Oct 9 18:32:14 2020     In reply to: 15619     Reply to this: 15626
Author: anchal 
Type: Summary 
Category: ALS 
Subject: ALS noise budget update - Updated AUX PDH Loop values 

The only two PZT Phase modulation transfer function measurements I could find are 40m/15206 and 40m/12077. Both these measurements were made to find a good modulation frequency and do not go below 50 kHz. So I don't think these will help us. We'll have to do a frequency transfer function measurement at lower frequencies.
I'm still looking for ALS PDH loop measurements to verify the model. I found this 40m/15059 but it is only near the UGF. The UGF measured here though looks very similar to the model prediction. A bit older measurement in 2017 was this 40m/13238 where I assume by ALS OLTF gautum meant the green laser PDH OLTF. It had similar UGF but the model I have has more phase lag, probably because of a 31.5 kHz pole which comes at U7 through the input low pass coupling through R28, C20 and R29 (See D1400293)

If the green laser is not being used, can I go and take some of these measurements myself?

ELOG V3.1.3-